OpenStudy (anonymous):

Sarah's craft project uses pieces of yarn that are 1/8 yard long. She has a piece of yarn that is 3 yards long. How many 1/8 yard pieces can she cut and still have 1 1/4 yards left?

OpenStudy (anonymous):

Are you supposed to write an equation to model this, then solve it? Or just figure it out any way you can?

OpenStudy (anonymous):

I need to draw a model.

OpenStudy (anonymous):

Then write my equation

OpenStudy (anonymous):

I'm not sure what you're looking for with the model, but I can help with the equation. If x represents the number of small pieces, then you need to multiply x by 1/8 since that's how long each of the small pieces are. Then if you add 1 1/4 to that product, you should get your total of 3 yards of yarn: \[\frac{ 1 }{ 8 }x + 1\frac{ 1 }{ 4 }=3\] If you also have to SOLVE that equation, I'd suggest either changing those fractions to decimals or multiplying by something to get rid of the fractions first.

OpenStudy (anonymous):

Thank you, I wlll try to work on it!

OpenStudy (anonymous):

Let me know if you need help! Good luck!

OpenStudy (anonymous):

Ok